# Simple Scaling Laws for Energy Correlators in Nuclear Matter

### v1

@@ Abstract
The remarkable aspect of the light-ray OPE is its prediction of the functional form of the leading nuclear modifications, which are quadratic in $\theta$, with the coefficient given by the expectation value of a twist-4 operator in the nuclear state. This matrix element can either be fitted to data or, enabling the determination of its dependence on the state’s properties. We will show that, for jet radii used in experiments, including up to the twist-4 contribution provides an excellent description of the data.
@@
Since the expectation values of the twist-4 operators $\langle\Psi_{\rm N}|\mathbb{O}_{\tau=4}^{[J=3]}|\Psi_{\rm N}\rangle$ in different states are unknown, we treat them as parameters to be fitted to the data in the lowest jet $p_{T}$ bin. A common feature of twist expansions is that matrix elements are accompanied by factors of $p_{T}^{-\tau}$ in the large-$p_{T}$ limit, fixed classically by dimensional analysis Luo et al. (1994a, b, 1992a); Qiu and Sterman (1992); Luo et al. (1992b, 1991); Kastella et al. (1989a, b). We thus determine the twist-4 matrix elements in higher $p_{T}$ bins by scaling the fitted result from the lowest bin according to this power law. To describe the data we must also incorporate the anomalous dimensions of the twist-two operators arising from their renormalization Dixon et al. (2019). We treat the twist-4 operators as classical, omitting extra parameters for their presently unknown anomalous dimensions. Incorporating these dimensions will be an interesting direction for future work, motivating further exploration of their structure Braun et al. (2001a, b); Derkachov et al. (2000); Braun et al. (2000, 1999); Homrich et al. (2024); Ekhammar et al. (2024); Belitsky (2000, 1999).
In Fig. 1, we show the Pb-Pb/p-p EEC ratio corrected by the unbiasing function $C_{2}$ Andres et al. (2024c); Andres and Holguin (2024) compared to the power-law scaling in (6), effectively of the form $\sim 1+a\theta^{2}$. The details of our fitting procedure are provided in the Supplemental Material, along with results for other $p_{T}$ and centrality bins. We see that the combination of the first two leading terms in the OPE (black curve) describe the data very well, with no free parameters required to describe the wide-angle enhancement in the middle and bottom panels of Fig. 1. It is genuinely remarkable that such a simple power-law can be identified in these complex collisions, as it also allows the measurement of the effective twist-4 matrix element characterizing it, which is sensitive to the properties of the medium. Additionally, this scaling reduces the calculation of the EEC to determining this matrix element, rather than determining a full function of $\theta$.
@@
In this Letter we have shown that the light-ray OPE provides an organizing principle for understanding the modification of energy correlators in nuclear collisions, by enabling a systematic expansion in the angle $\theta$, with the coefficients of distinct power-law scalings associated with matrix elements of light-ray operators. We showed that the leading nuclear modification is associated with enhanced twist-4 light-ray operators, and that by incorporating just these operators, we are able to predict the $\theta$ dependence within the range of angles relevant for jet substructure studies. Our prediction beautifully describes both the recent CMS data for the EEC in A-A collisions, and the ALICE data for the EEC in p-A.
@@
${\rm N/pp}=\mathcal{M}_{2}\frac{\theta^{\gamma_{N}}+\mathcal{M}_{4}~{}\theta^{2}}{\theta^{\gamma}}\,,$ (14)
@@
In summary, we fit both the A-A and p-A data with the following ansatz
${\rm N/pp}=\mathcal{M}_{2}\frac{\theta^{\gamma_{N}}+\left(\frac{p_{T,0}}{p_{T}}\right)^{2}\mathcal{M}_{4}(p_{T,0})~{}\theta^{2}}{\theta^{\gamma}}\,,$ (15)
where $\mathcal{M}_{4}(p_{T,0})$ is fitted to a particular reference bin in $p_{T,0}$. Therefore, outside the reference $p_{T}$ bin, there are no unconstrained parameters. In the future, we can improve this analysis by directly computing the q/g fractions and twist-4 anomalous dimensions. In this case, the only remaining parameter will be $\mathcal{M}_{4}(p_{T,0})$ which is strongly dependent on the properties of the process. $\mathcal{M}_{4}(p_{T,0})$ will therefore provide a highly sensitive probe of the nuclear medium.
@@
In Fig. 4 we show the OPE fitted to the unbiased CMS Pb-Pb EEC measurement CMS (2024); Andres et al. (2024c); Andres and Holguin (2024) across 4 bins in $p_{T}$ and 4 bins in centrality. We use the 120-140 GeV $p_{T}$ bin (the top row) to find $\mathcal{M}_{4}(120\mbox{-}140~{}{\rm GeV})$ for each centrality class. The functional form of the wide angle enhancement is then entirely fixed in rows 2-4. It is clear that excellent agreement with the experimental data is found. Comparing to the analytic calculation with twist-2 leading-log resummation (which will be discussed in the following paragraph) we found that anomalous dimensions are consistent with a fractionally higher quark fraction in Pb-Pb than in p-p of $\sim 10\%$.
In Fig. 2 of the main body of the letter we apply the OPE ansatz to the p-A measurement. As was discussed in the main body of this letter, in this case the effective anomalous dimensions ($\gamma_{pA}$ and $\gamma$) play a large roll in the lowest $p_{T}$ bin. Using leading resummation for the left-hand-side of eq. (13), one can solve for the quark fraction in p-A in terms of the p-p quark fraction which results in the measured values of $\gamma_{pA}$ and $\gamma$. The relevant leading-log formulae are derived in full in the appendices of Andres et al. (2023b). The relevant extracted fit parameters were $\gamma_{pA}=0.34$ and $\gamma=0.29$ which is consistent with $N^{pA}_{q}\approx-0.05+0.85N^{pp}_{q}$, where we normalize the q/g fractions so that $N_{q}+N_{g}=1$. Therefore, assuming an even q/g fraction in p-p, one finds an approximate p-A q/g fraction of $30/70$ consistent with expectations from nuclear PDFs Abdul Khalek et al. (2022b). An important observation in the p-Pb fit is that the $p_{T}$ suppression of the twist-4 contribution has a dominant effect in the higher $p_{T}$ bins. This justifies the absence of any enhancement in p-Pb/p-p ratio for higher jet $p_{T}$.

### v2

@@ Abstract
The remarkable aspect of the light-ray OPE is its prediction of the functional form of the leading nuclear modifications, which are quadratic in $\theta$, with the coefficient given by the expectation value of a twist-4 operator in the nuclear state. This matrix element can either be fitted to data or analytically computed, enabling the determination of its dependence on the state’s properties. We will show that, for jet radii used in experiments, including up to the twist-4 contribution provides an excellent description of the data.
@@
Since the expectation values of the twist-4 operators $\langle\Psi_{\rm N}|\mathbb{O}_{\tau=4}^{[J=3]}|\Psi_{\rm N}\rangle$ in different states are unknown, we treat them as parameters to be fitted to the data in the lowest jet $p_{T}$ bin. A common feature of twist expansions is that matrix elements are accompanied by factors of $p_{T}^{-\tau}$ in the large-$p_{T}$ limit, fixed classically by dimensional analysis Luo et al. (1994a, b, 1992a); Qiu and Sterman (1992); Luo et al. (1992b, 1991); Kastella et al. (1989a, b). We thus determine the twist-4 matrix elements in higher $p_{T}$ bins by scaling the fitted result from the lowest bin according to this power law. To describe the data, we must also incorporate the anomalous dimensions of the twist-two operators arising from their renormalization Dixon et al. (2019). Twist-4 operators are treated as classical, omitting extra parameters for their presently unknown anomalous dimensions. Including these dimensions will be an interesting future direction, motivating further exploration of their structure Braun et al. (2001a, b); Derkachov et al. (2000); Braun et al. (2000, 1999); Homrich et al. (2024); Ekhammar et al. (2024); Belitsky (2000, 1999).
In Fig. 1, we show the Pb-Pb/p-p EEC ratio corrected by the unbiasing function $C_{2}$ Andres et al. (2024c); Andres and Holguin (2024) compared to the power-law scaling in (6), effectively of the form $\sim 1+a\theta^{2}$. The details of our fitting procedure are provided in the Supplemental Material, along with results for other $p_{T}$ and centrality bins. We see that the combination of the first two leading terms in the OPE (black curve) describes the data very well, with no free parameters required to describe the wide-angle enhancement in the middle and bottom panels of Fig. 1. It is genuinely remarkable that such a simple power-law can be identified in these complex collisions, as it also allows the measurement of the effective twist-4 matrix element characterizing it, which is sensitive to the properties of the medium. Additionally, this scaling reduces the calculation of the EEC to determining this matrix element, rather than determining a full function of $\theta$.
@@
Energy correlators can be formulated directly in terms of light-ray operators, in contrast to traditional jet substructure observables. In this Letter we have shown that the light-ray OPE provides an organizing principle for understanding the modification of energy correlators in nuclear collisions, by enabling a systematic expansion in the angle $\theta$, with the coefficients of distinct power-law scalings associated with matrix elements of light-ray operators. We showed that the leading nuclear modification is associated with enhanced twist-4 light-ray operators, and that by incorporating just these operators, we are able to predict the $\theta$ dependence within the range of angles relevant for jet substructure studies. Our prediction beautifully describes both the recent CMS data for the EEC in A-A collisions, and the ALICE data for the EEC in p-A.
@@
${\rm N/pp}=\frac{\mathcal{M}_{2}~{}\theta^{\gamma_{N}}+\mathcal{M}_{4}~{}\theta^{2}}{\theta^{\gamma}}\,,$ (14)
@@
In summary, we fit both the A-A/p-p and p-A/p-p data with the following ansatz
${\rm N/pp}=\frac{\mathcal{M}_{2}~{}\theta^{\gamma_{N}}+\left(\frac{p_{T,0}}{p_{T}}\right)^{2}\mathcal{M}_{4}(p_{T,0})~{}\theta^{2}}{\theta^{\gamma}}\,,$ (15)
where we have made explicit the $p_{T}$-dependence of the twist-4 matrix element relative to the twist-2 term. Here, $\mathcal{M}_{4}(p_{T,0})$ is fitted to a particular reference bin in $p_{T}$, $p_{T,0}$, for p-A/p-p and to a particular reference bin in $p_{T}$, $p_{T,0}$, for each centrality class in A-A/p-p data. The reference value of $p_{T,0}$ is chosen to be the lowest $p_{T}$ bin for which data are available, both for the p-A/p-p and A-A/p-p fits. Consequently, in p-A, $\mathcal{M}_{4}(p_{T,0})$ remains constant across all $p_{T}$ bins. This means that the modification (if any) obtained in the two higher $p_{T}$ bins in p-A, relative to p-p, is a predicted outcome of the $p_{T}$-dependence of the twist of the matrix elements. Similarly, in the A-A/p-p fits, $\mathcal{M}_{4}$ remains constant across all $p_{T}$ bins within a given centrality class. As a result, the size of the A-A enhancement relative to p-p is a prediction for all $p_{T}$ bins within each centrality class, except for the bin used to determine $\mathcal{M}_{4}$. Additionally, since $\gamma$ is determined by the p-p baseline and should therefore be independent of centrality, it is allowed to vary only as a function of $p_{T}$. In the A-A/p-p fits, $\gamma$ is thus fixed for each $p_{T}$ bin across all centralities to the value obtained from the most central class. In the future, this analysis could be improved by directly computing the q/g fractions and twist-4 anomalous dimensions. With these additional inputs, the only remaining parameter would be $\mathcal{M}_{4}(p_{T,0})$, which is strongly dependent on the properties of the process. As such, $\mathcal{M}_{4}(p_{T,0})$ would provide a highly sensitive probe of the nuclear medium.
@@
In Fig. 4 we show the OPE fitted to the unbiased CMS Pb-Pb/p-p EEC measurement CMS (2024); Andres et al. (2024c); Andres and Holguin (2024) across 4 bins in $p_{T}$ and 4 bins in centrality. We use the 120-140 GeV $p_{T}$ bin (the top row) to find $\mathcal{M}_{4}(120\mbox{-}140\,{\rm GeV})$ for each centrality class. The functional form of the wide angle enhancement is then entirely fixed in rows 2-4. It is clear that excellent agreement with the experimental data is found. Comparing to the analytic calculation with twist-2 leading-log resummation (which will be discussed in the following paragraph) we found that anomalous dimensions are consistent with a fractionally higher quark fraction in Pb-Pb than in p-p of $\sim 10\%$. We note that setting $\mathcal{M}_{4}(120\mbox{-}140\,{\rm GeV})=0$ fails to describe the data, indicating that twist-4 contributions are essential.
In Fig. 2 of the main body of the Letter we apply the OPE ansatz to the p-A/p-p ratio. As was discussed in the main body of this Letter, in this case the effective anomalous dimensions ($\gamma_{pA}$ and $\gamma$) play a large roll in the lowest $p_{T}$ bin. Using leading resummation for the left-hand-side of (13), one can solve for the quark fraction in p-A in terms of the p-p quark fraction which results in the measured values of $\gamma_{pA}$ and $\gamma$. The relevant twist-2 leading-log formulae are derived in full in the appendices of Andres et al. (2023b). The relevant extracted fit parameters were $\gamma_{pA}=0.34$ and $\gamma=0.29$ are consistent with $N^{pA}_{q}\approx-0.05+0.85N^{pp}_{q}$, where we normalize the q/g fractions so that $N_{q}+N_{g}=1$. Therefore, assuming an even q/g fraction in p-p, one finds an approximate p-Pb q/g fraction of $30/70$ consistent with expectations from nuclear PDFs Abdul Khalek et al. (2022b). An important observation in the p-Pb/p-p fit is that the $p_{T}$-suppression of the twist-4 contribution has a dominant effect in the higher $p_{T}$ bins. This justifies the absence of any enhancement in p-Pb/p-p ratio for higher jet $p_{T}$.